A company uses H2 gas to generate electricity. They studied a smaller scale model

reaction to figure out the maximum amount of H2 gas that they can produce. For this

purpose, they used an alloy consisting of 95% Al by mass. The alloy has a density of 3.12

g/cm3 and they took a 0.875 cm3 small piece of alloy to react with 200 mL of 1.2 M solution

of HCl based on the given Al specific reaction. What is the maximum amount of H2 gas

that can be produced? What are the limiting and remaining reagents? How can they

increase the amount of H2 gas that they obtained?


Al + HCl → AlCl3 + H2 (not balanced)


1
Expert's answer
2021-12-14T12:59:01-0500

On way of working this type of problem is to assume that one starts with an alloy mass of 1 kg = 1000 g.

Then there would be 0.70 x 1000 = 700 g oif Al and .30 x 1000 = 300 g Mg.

 

Since the Molar Mass of Al is 26.982 g /mole , there are 700/ 26.982 = 25.94 molles of Al, similarly there are

 300 / 24.303 = 12.34 moles of Mg  (Molar mass of Mg is 24.305)

Thus the mole fraction of Al ( f Al ) is 25.94/(25.94 + 12.34) = 0.6776

and the mole fraction of Mg (f Mg ) is  12.34 /(25.98 + 12.34) = 0.3224

 

The density of alloy will then be 0.6776 x 2700 + 0.3324 x 1740 = 2390 kg / m^3